Calculate GDP using the Income and Expenditure Approach. (Both approach should be the same answer). Show all working for both Approaches

(all figures are in billions of dollars):


Item Amount ($)


Government purchase of goods and services 1,721.6

Exports 1,096.3

Receipts of factor income from the rest of the world 382.7

Depreciation (consumption of fixed capital) 990.8

Net fixed Investments 688.2

Corporate income taxes 265.2

Consumption expenditures 6,739.4

Indirect business taxes 664.6

Imports 1,475.8

Payments of factor income to the rest of the world 343.7

Inventory change 56.5

Social security contributions 702.7

Undistributed corporate profits (retained earnings) 130.3

Government transfer and interest payments 1,366.3

Personal interest payments 286.2

Personal taxes 1,235.7


1
Expert's answer
2021-11-16T11:36:13-0500

Income approach:


=Receipts of factor income from the rest of the world +Depreciation +corporate income taxes+ indirect business taxes+ payments of factor income to the rest of the world+ social security contributions+ undistributed corporate taxes+ government transfer and interest payments+ personal interest payments+ personal taxes


"=382.7+ 990.8+ 265.2+ 664.6+ 343.7+ 702.7+ 130.3+ 1366.3+ 1235.7+286.2=\$6368.2"


Expenditure approach;


=government purchase of goods and services+ net fixed investments+ consumption expenditure+ inventory change+(exports-imports)


"=1721.6+688.2+6739.4+56.5+(1096.3-1475.8)" ="\$8826.2"